OpenStudy (anonymous):

Find the product. y 4 · y 3

OpenStudy (sweetburger):

what is y·y·y·y·y·y·y =?

OpenStudy (anonymous):

12y^2

OpenStudy (anonymous):

Is that what you are looking for?

OpenStudy (anonymous):

y*4*y*3 =12*y^2 =12y^2

OpenStudy (sweetburger):

\[(y^4)(y^3) = ?\] is this your question?

OpenStudy (anonymous):

yes

OpenStudy (butterflydreamer):

note: When multiplying the same base, we ADD the exponents (power) So use this exponent rule :)